Find a quadratic equation with roots -1 + 4i and -1 - 4i.
Answer by stanbon(75887)

Find a quadratic equation with roots -1 + 4i and -1 - 4i
--------------------------
f(x) = (x-(-1+4i))(x-(-1-4i))
Rearrange:
f(x) = [(x+1)-4i][(x+1)+4i]
f(x) = (x+1)^2 - (4i)^2
f(x) = x^2 + 2x + 1 + 16
f(x) = x^2 + 2x + 17
=======================
Cheers,
Stan H.